OpenStudy (anonymous):

A body was found at 10 a.m. outdoors on a day when the temperature was 40oF. The medical examiner found the temperature of the body to be 80oF. What was the approximate time of death? Use Newton's law of cooling, with k = 0.1947. \[T(t)=T _{A}+(T _{o}-T _{A)}e^{-kt}\] A. 5 am B. 8 am C. 9:45 am D. 9:30 am

OpenStudy (anonymous):

work with the differences between the outdoor temperatures and the body temp at 10 the difference is \(80-40=40\) and presumably at the time of death it was \(98.6\) so the difference was \(98.6-40=58.6\) set \[\large 58.6=40e^{1947 t}\] and solve for \(t\)

OpenStudy (anonymous):

sorry typo on the last line, should be \[\large 58.6=40e^{-.1947 t}\]

OpenStudy (anonymous):

you know how to solve this?

OpenStudy (anonymous):

no i dont actually, could you help me out?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

mhhm

OpenStudy (anonymous):

step one, divide both sides by \(40\) and get \[\frac{58.9}{40}=e^{-.1947t}\]

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

quick question when you add or subtract an inequality do you flip the sign?

OpenStudy (anonymous):

damn another typo, i meant \[\frac{58.6}{40}=e^{-1947t}\] or \[1.465=e^{-1947t}\]

OpenStudy (anonymous):

I got 1.4725 when I divided by 40, but what do I do with the e^-.1947t ?

OpenStudy (anonymous):

no you only flip the inequality sign when you multiply or divide. @Karina1

OpenStudy (anonymous):

yeah i made a typo there, it should be \(58.6\div 40=1.465\)

OpenStudy (anonymous):

oh okay, and what do I do for the next step?

OpenStudy (anonymous):

is the multiplication (does the) have a impact on the flip, like i know you flip it, bury I'm confused on that part

OpenStudy (anonymous):

so you have \[e^{-.1947t}=1.465\] in equivalent logarithmic form you have \[-.1947t=\ln(1.465)\]

OpenStudy (anonymous):

then divide to get \[t=-\frac{\ln(1.465)}{.1947}\]

OpenStudy (anonymous):

thanks @ali1029

OpenStudy (anonymous):

i get \(-1.96\) so about 2 hours before 10

OpenStudy (anonymous):

I divided that and got -1.96, is that right?

OpenStudy (anonymous):

so it would be 8 am?

OpenStudy (anonymous):

yes that is what i get too

OpenStudy (anonymous):

thank you so much! :)